A cement manufacturer buys limestone for $3 per pound and a special shale mix for $5 per pound. The two ingredients are mixed into a 100-pound compound that cost $4.25 per pound. How much limestone and how much of the shale mix should be added into the compound?
Answer by checkley71(8403)

4.25*100=3.00*X+5.00(100-X)
425=3X+500-5X
2X=500-425
2X=75
X=75/2
X=37.5 POUNDS OF LIMESTONE
100-37.5=62.5 POUNDS OF SHALE
PROOF
425=3*37.5+5*62.5
425=112.5+312.5
425=425